# iSeg: An Iterative Refinement-based Framework for Training-free Segmentation

### v2

@@ Abstract
Stable diffusion has demonstrated strong image synthesis ability to given text descriptions, suggesting it to contain strong semantic clue for grouping objects. Inspired by this, researchers have explored employing stable diffusion for training-free segmentation. Most existing approaches either simply employ cross-attention map or refine it by self-attention map, to generate segmentation masks. We believe that iterative refinement with self-attention map would lead to better results. However, we empirically demonstrate that such a refinement is sub-optimal likely due to the self-attention map containing irrelevant global information which hampers accurately refining cross-attention map with multiple iterations. To address this, we propose an iterative refinement framework for training-free segmentation, named iSeg, having an entropy-reduced self-attention module which utilizes a gradient descent scheme to reduce the entropy of self-attention map, thereby suppressing the weak responses corresponding to irrelevant global information. Leveraging the entropy-reduced self-attention module, our iSeg stably improves refined cross-attention map with iterative refinement. Further, we design a category-enhanced cross-attention module to generate accurate cross-attention map, providing a better initial input for iterative refinement. Extensive experiments across different datasets and diverse segmentation tasks (weakly-supervised semantic segmentation, open-vocabulary semantic segmentation, unsupervised segmentation, and mask generation on synthetic dataset) reveal the merits of proposed contributions, leading to promising performance on diverse segmentation tasks. For unsupervised semantic segmentation on Cityscapes, our iSeg achieves an absolute gain of in terms of mIoU compared to the best existing training-free approach in literature. Moreover, our proposed iSeg can support segmentation with different kind of images and interactions. Our project page is publicly released at https://linsun449.github.io/iSeg/.
1 Introduction
@@
Semantic segmentation [1, 2, 3] aims to differentiate the pixels into different semantic groups. In past decade, fully-supervised semantic segmentation has achieved significant progress where the models are trained on large-scale datasets with pixel-level annotations and fixed categories. This fully-supervised setting hinders zero-shot segmentation ability on unseen categories. Further, it is expensive to collect and annotate large-scale dataset with numerous categories [4] for model training. One straightforward solution is to develop a universal and training-free framework that exhibits strong generalization capabilities across arbitrary categories without any training on segmentation datasets.
@@
Recent work of [5] introduces a vision-language model, CLIP, that learns the connection between text and images through training on large-scale image-text pairs, showing promising generalization capabilities in recognizing arbitrary categories. Inspired by this, several works [6, 7, 8, 9] explore employing the strong zero-shot ability of CLIP for semantic segmentation. Some works focus on directly utilizing the learned text-image mapping from pre-trained CLIP. For instance, OVSeg [6] first utilizes mask proposal generator to extract some class-agnostic masks, and then employs CLIP to classify cropped and masked images for open-vocabulary semantic segmentation. Some works employ text supervision similar to CLIP. For instance, GroupViT [8] introduces a hierarchical grouping vision transformer with text supervision, which is able to group different semantic regions without pixel-level annotations. These methods improve the generalization ability of semantic segmentation or reduce the reliance on pixel-level annotations in some degree. However, CLIP focuses on image-level information, limiting its capability in pixel-level tasks, especially under training-free setting.
@@
Recently, researchers started to employ text-to-image stable diffusion model for segmentation. Stable diffusion [10] was initially designed to perform high-quality image synthesis. With the pre-training on large-scale text-image paired data, stable diffusion is able to generate high-quality images with reasonable structure according to given text descriptions. Inspired by this, researchers believed that stable diffusion contains rich semantic clue to group objects, and proposed various diffusion-based segmentation approaches. Some approaches [11, 12, 13, 14] treat stable diffusion as feature extractor, and re-train it on segmentation dataset. In contrast, other approaches [15, 16, 17] explore training-free framework by utilizing self-attention map and cross-attention map in stable diffusion. For instance, DiffSeg [15] iteratively merges self-attention maps at different positions and employs non-maximum suppression to assign the pixels into different objects. DiffuMask [18] generates initial object masks of synthetic images using cross-attention maps. DiffSegmenter [16] refines the cross-attention map with self-attention map for improved segmentation. These approaches present an initial attempt on employing self-attention and cross-attention maps for training-free segmentation tasks. However, we argue that it is still an open problem about how to fully harness self-attention and cross-attention maps for segmentation tasks.
@@
To fully exploit the potential ability of cross-attention and self-attention maps in stable diffusion for training-free segmentation, we conduct a preliminary experiment, where we iteratively refine the cross-attention map with self-attention map as in Fig. 1. Given a text prompt (a) and an image (b), we employ stable diffusion to extract cross-attention map (c) of category bird. Subsequently, we perform a naive iteration strategy to refine the cross-attention map by iteratively multiplying self-attention map. After several iterations, we observe that the refined cross-attention map associated with category bird gradually has strong response on other regions, such as person and sandwich in (d). We argue that this degradation in refined cross-attention map is caused by self-attention map that aggregates global information from irrelevant regions after multiple iterations. To address this issue, we propose an iterative refinement framework by introducing an entropy-reduced self-attention (Ent-Self) module to impede the information aggregation from irrelevant regions. As in (e), our Ent-Self module is able to generate a stable and accurate cross-attention map of bird after multiple iterations. We further give a quantitative comparison of generated pseudo mask for weakly-supervised semantic segmentation in (f). Compared to naive iteration strategy, our Ent-Self module exhibits improved performance with the increment of iterations. In addition, we design a simple category-enhanced cross-attention (Cat-Cross) module to generate more accurate cross-attention map as the input fed to iterative refinement. The Cat-Cross module employs a weighting strategy to enhance the text embeddings of given category while suppressing other words. We perform the experiments on various datasets and segmentation tasks, which demonstrate method efficacy.
2 Related works
@@
Semantic segmentation. Initially, researchers mainly investigated CNN-based networks. One of the pioneering approaches is FCN [19], which utilizes convolutional neural networks with skip-layer connections. Subsequently, numerous variants were developed, incorporating innovations such as spatial pyramid structures [1, 20], multi-layer feature fusion [21, 22, 23], and non-local enhancements [24, 25, 26]. With the success of vision transformer in image classification [27, 28], researchers began to explore transformer-based approaches. Some studies [29, 30, 3] focus on harnessing the powerful representational capabilities of transformer by employing it as backbone. Additionally, other researches [31, 32] conceptualize semantic segmentation as a set prediction task, utilizing the transformer as decoder.
@@
Besides, researchers developed various semantic segmentation paradigms. Weakly-supervised semantic segmentation [33, 34, 35] solely relies on image-level category annotations to train models. For this task, many researchers employed CAM [36] to generate pseudo pixel-level annotations and then trained models with these annotations. Open-vocabulary semantic segmentation [37, 38, 8] performs semantic segmentation on unseen categories. In open-vocabulary semantic segmentation, researchers mainly explored to utilize CLIP [5] to segment unseen categories. Unsupervised segmentation [39, 40, 41] aims to group the pixels belonging to the same objects without using any annotations.
@@
Diffusion models for segmentation. Diffusion models [42] have exhibited strong image synthesis ability across various text prompts, indicating their strong semantic clue for image segmentation. Inspired by this, researchers started to employ the pre-trained diffusion models for segmentation. For instance, VPD [13] views stable diffusion model as feature extractor, and fine-tunes them for semantic segmentation tasks. TADP [43] introduces image-aligned text prompts to condition diffusion model for segmentation. SLIME [44] adopts stable diffusion model with learnable text embeddings for few-shot segmentation. Additionally, some researchers [45] adopted diffusion framework for predicting accurate segmentation maps from noisy data.
@@
Recently, some researchers explored training-free diffusion models for segmentation. One of the key idea is to employ self-attention map and cross-attention map from pre-trained diffusion model to generate masks of semantic categories. For instance, DiffuMask [18] first employs diffusion model to generate synthetic images, and calculates the averaged cross-attention map as the initial object masks. DiffSeg [15] explores grouping self-attention maps at different positions into distinct objects based on their similarities, followed by non-maximum suppression to assign pixels to individual objects. OVDiff [46] generates synthetic images by diffusion model, and uses a feature extractor to generate the foreground and background prototypes of different categories for open-vocabulary segmentation. Both Diffsegmenter [16] and T2M [17] refine cross-attention maps of different categories with self-attention maps. To improve cross-attention map, Diffsegmentor and T2M employ additional BLIP [47] and CLIP to carefully design text prompts, which are relatively complex in some degree. Though these aforementioned approaches exploit self-attention and cross-attention maps for training-free segmentation, we argue that these approaches do not fully utilize the semantic clue existing in self-attention and cross-attention maps. To address this issue, we propose an iterative training-free segmentation framework using pre-trained stable diffusion model.
@@ 3.1 Preliminaries
Stable diffusion model. Text-to-image (T2I) diffusion model is a type of text-conditioned diffusion models [42], such as DALL-E [48], GLIDE [49], Imagen [50], and stable diffusion [10]. As one of representative T2I models, stable diffusion generates high-quality images with text prompts, which is trained on large-scale text-image paired dataset (e.g., LAION-5B [51]). It comprises three components: (1) A pre-trained VAE [52] encoder and decoder for image compression and restoration. (2) A text encoder that encodes text descriptions into latent semantic space. (3) A transformer-based denoising UNet [53] that recovers images from noisy data according to text descriptions.
Attention modules. Denoising UNet in stable diffusion integrates self-attention and cross-attention mechanisms. The self-attention models the relation between different spatial positions of visual features, while the cross-attention models the relation between visual features and text embeddings. The attention map in both self-attention and cross-attention operations can be calculated by the query and the key as , where is normalization factor. Assuming that the visual features has the spatial size of , and the text embeddings has the length of . In self-attention operation, the query and key are both generated by visual features via linear layers. The generated self-attention map is denoted as , which can reflect spatial similarities of different positions within image. In cross-attention operation, the query and key are respectively generated by visual features and text embeddings via linear layers. The generated cross-attention map is denoted as , capturing the similarities between image and text.
To achieve training-free segmentation, a straightforward way is setting the text prompts based on given category names, such as ‘A photograph of cat and dog’, then performing the forward process of stable diffusion and finally generating the corresponding cross-attention map between visual features and text embeddings. Afterwards, we can predict segmentation map based on the obtained cross-attention map associated with corresponding category name. Instead of directly using cross-attention map, some researchers [44] proposed to generate refined cross-attention map with self-attention map, which can be written as
@@ 3.2 Motivation
Problem. As mentioned above, self-attention map can be used to refine the cross-attention map. Therefore, a natural question is that, can we improve cross-attention map iteratively with self-attention map, formulated as , where is the index of iterations. We initially believe that this iterative refinement would yield improved segmentation performance. However, as in Fig. 1(f), it presents performance degradation after several iterations. To better analyze this question, we perform the experiment in Fig. 2. Given the input image (a) and cross-attention map (b), we present the refined cross-attention map at different iterations in (c). It can be observed that the refinement through multiple iterations can not guarantee to generate better refined cross-attention map. We divide the refined results into three situations. The first situation is that the cross-attention map becomes better with the increasing iterations, such as cross-attention map of dog in first row. The second one is the cross-attention map becomes larger, such as the cross-attention map of sheep in the second row. The last one is that the cross-attention map shifts to other regions, exemplified by the cross-attention map of dog showing strong response on the person in third row.
Analysis. Ideally, we argue the refinement with multiple iterations at least presents a stable cross-attention map around corresponding category if the self-attention map is relatively accurate. We further perform a simple experiment by refining cross-attention map with ground-truth self-attention map in Fig. 2(d). The ground-truth self-attention map is generated by self-correlation of ground-truth masks, which is binary and exclusively has responses within object regions. By using ground-truth self-attention map, the refined cross-attention map becomes stable and accurate after only one iteration. This observation demonstrates that an accurate self-attention map is crucial for accurately refining cross-attention map. As shown in (e), we compare the predicted self-attention map (middle) and ground-truth self-attention map (right) at a given point . Compared to ground-truth self-attention map, predicted self-attention map is relatively noisy and has some weak responses at irrelevant regions. As a result, the refinement with predicted self-attention map aggregates irrelevant global information into the refined cross-attention map. Therefore, we argue that it is important to reduce irrelevant global information within self-attention map for iterative refinement.

### v3

@@ Abstract
Stable diffusion has demonstrated strong image synthesis ability to given text descriptions, suggesting it to contain strong semantic clue for grouping objects. The researchers have explored employing stable diffusion for training-free segmentation. Most existing approaches refine cross-attention map by self-attention map once, demonstrating that self-attention map contains useful semantic information to improve segmentation. To fully utilize self-attention map, we present a deep experimental analysis on iteratively refining cross-attention map with self-attention map, and propose an effective iterative refinement framework for training-free segmentation, named iSeg. The proposed iSeg introduces an entropy-reduced self-attention module that utilizes a gradient descent scheme to reduce the entropy of self-attention map, thereby suppressing the weak responses corresponding to irrelevant global information. Leveraging the entropy-reduced self-attention module, our iSeg stably improves refined cross-attention map with iterative refinement. Further, we design a category-enhanced cross-attention module to generate accurate cross-attention map, providing a better initial input for iterative refinement. Extensive experiments across different datasets and diverse segmentation tasks (weakly-supervised semantic segmentation, open-vocabulary semantic segmentation, unsupervised segmentation, and mask generation on synthetic dataset) reveal the merits of proposed contributions, leading to promising performance on diverse segmentation tasks. For unsupervised semantic segmentation on Cityscapes, our iSeg achieves an absolute gain of in terms of mIoU compared to the best existing training-free approach in literature. Moreover, our proposed iSeg can support segmentation with different kinds of images and interactions. The project is available at https://linsun449.github.io/iSeg.
Index Terms:
Training-free segmentation, stable diffusion, iterative refinement, entropy-reduced self-attention, category-enhanced cross-attention.1 Introduction
@@
Semantic segmentation [10, 62, 68] aims to differentiate the pixels into different semantic groups. In past decade, fully-supervised semantic segmentation has achieved significant progress where the models are trained on large-scale dataset with pixel-level annotations and fixed categories. This fully-supervised setting hinders zero-shot segmentation ability on unseen categories. Further, it is expensive to collect and annotate large-scale dataset with numerous categories [35] for model training. One straightforward solution is to develop a universal and training-free framework that exhibits strong generalization capabilities across arbitrary categories without any training on segmentation datasets.
@@
Recent work of [48] introduces a vision-language model, CLIP, that learns the connection between text and image through training on large-scale image-text pairs, showing promising generalization capabilities in recognizing arbitrary categories. Inspired by this, several works [34, 37, 70, 69] explore employing the strong zero-shot ability of CLIP for semantic segmentation. Some works focus on directly utilizing the learned text-image mapping from pre-trained CLIP. For instance, OVSeg [34] first utilizes mask proposal generator to extract some class-agnostic masks, and then employs CLIP to classify the cropped and masked images for open-vocabulary semantic segmentation. Some works employ text supervision similar to CLIP. For instance, GroupViT [70] introduces a hierarchical grouping vision transformer with text supervision, which is able to group different semantic regions without pixel-level annotations. These methods are able to improve the generalization ability of semantic segmentation or reduce the reliance on pixel-level annotations in some degree. However, CLIP focuses on image-level information, limiting its capability in pixel-level segmentation tasks, especially under training-free setting.
@@
Recently, researchers started to employ text-to-image stable diffusion model for segmentation. Stable diffusion [52] was initially designed to perform high-quality image synthesis. With the pre-training on large-scale text-image paired data, stable diffusion is able to generate high-quality images with reasonable structure according to given text descriptions. Inspired by this, researchers believed that stable diffusion contains rich semantic clue to group objects, and proposed various diffusion-based segmentation approaches. Some approaches [64, 4, 77, 60] treat stable diffusion as feature extractor, and re-train it on segmentation dataset. In contrast, other approaches [59, 63, 67] explore training-free framework by utilizing self-attention map and cross-attention map in stable diffusion. For instance, DiffSeg [59] iteratively merges self-attention maps at different positions and employs non-maximum suppression to assign the pixels into different objects. DiffuMask [65] generates initial object masks of synthetic images using cross-attention maps. DiffSegmenter [63] refines the cross-attention map with self-attention map for improved segmentation. These approaches present an initial attempt on employing self-attention and cross-attention maps for training-free segmentation tasks, demonstrating that self-attention and cross-attention maps contain useful semantic information to improve segmentation. Therefore, there is a natural question that, how to fully harness self-attention and cross-attention maps for segmentation tasks.
@@
In this paper, to fully exploit the potential ability of cross-attention and self-attention maps in pre-trained stable diffusion for training-free segmentation, we conduct a preliminary experiment, where we iteratively refine the cross-attention map with self-attention map as in Fig. 1. Given a text prompt (a) and an image (b), we employ pre-trained stable diffusion to extract cross-attention maps and select the cross-attention maps corresponding to different category names, such as people and bird, in (c). Subsequently, we perform a naive iteration strategy to refine the cross-attention maps by iteratively multiplying self-attention map. After several iterations, we observe that the refined cross-attention maps, associated with categories people and bird, gradually have strong response on other regions in (d). We argue that this degradation in refined cross-attention maps is caused by self-attention map that aggregates global information from irrelevant regions after multiple iterations.
To address this issue mentioned above, we propose an iterative refinement framework by introducing an entropy-reduced self-attention (Ent-Self) module to impede the information aggregation from irrelevant regions. As in (e), using our Ent-Self module is able to generate the stable and accurate cross-attention maps of person and bird after multiple iterations. We further give a quantitative comparison of generated pseudo mask for weakly-supervised semantic segmentation in (f). Compared to naive iteration strategy, using our Ent-Self module exhibits improved performance with the increment of iterations. In addition, we design a simple category-enhanced cross-attention (Cat-Cross) module to generate more accurate cross-attention maps corresponding to the category names, which are fed to iterative refinement. The Cat-Cross module employs a weighting strategy to enhance the text embeddings of given category names while suppressing the text embeddings of other words. We perform the experiments on various datasets and segmentation tasks, including weakly-supervised semantic segmentation, open-vocabulary semantic segmentation, unsupervised segmentation, and mask generation on synthetic dataset. Extensive experiments demonstrate the efficacy of our proposed method. The contributions and merits of our method can be summarized as
We give a deep analysis on simply refining cross-attention map using self-attention map with multiple iterations, and Analyze the reason why this naive iteration strategy presents worse performance.
We introduce an effective iterative refinement for improved segmentation using our proposed entropy-reduced self-attention module, which can avoid incorporate irrelevant global information within self-attention maps during multiple iterations.
A category-enhanced cross-attention map is proposed to enhance the text embeddings of given categories while suppressing other words.
The experiments are performed on various segmentation tasks, including weakly-supervised semantic segmentation, open-vocabulary semantic segmentation, unsupervised segmentation, and mask generation on synthetic dataset. Our proposed method can significantly improve segmentation accuracy.
Our proposed method can support the segmentation on different kinds of images or different interactions, such as point, line and box.
2 Related Work
In this section, we first give a review on semantic segmentation. Afterwards, we introduce the related works of using diffusion models for segmentation.
2.1 Semantic Segmentation
Semantic segmentation is one of classical computer vision tasks, which aims to group the pixels belonging to same semantic category together. In the past decade, semantic segmentation has achieved the great progress. The related methods can be mainly divided into CNN-based approaches and transformer-based approaches.
@@ 2.1 Semantic Segmentation
Initially, the researchers mainly investigated CNN-based networks. One of the pioneering approaches is FCN [40], which utilizes the convolutional neural networks with skip-layer connections for improved semantic segmentation. Subsequently, numerous variants were developed, incorporating innovations such as spatial pyramid structures [10, 9], multi-layer feature fusion [39, 3, 11], and non-local enhancements [19, 22, 75]. The spatial pyramid structure usually adopts multiple paralleled branches with different receptive fields to extract diverse context information. Multi-layer feature fusion employs skip-layer connection to fuse the feature maps of different layers. Non-local enhancement adopts the attention operation to integrate globally context information. With the success of vision transformer in image classification [17, 6], researchers began to explore transformer-based segmentation approaches. Some studies [74, 21, 68] focus on harnessing the powerful representational capabilities of transformer by employing it as backbone. Additionally, other researches [13, 12] conceptualize semantic segmentation as a set prediction task, utilizing the transformer as the decoder.
@@
These approaches mentioned above belong to fully-supervised semantic segmentation. Besides, researchers developed various semantic segmentation paradigms, such as weakly-supervised semantic segmentation, open-vocabulary semantic segmentation, and unsupervised segmentation. Compared to fully-supervised semantic segmentation, these tasks are more challenging. Weakly-supervised semantic segmentation [45, 47, 2] solely relies on image-level category annotations to train the models. For this task, many researchers employed CAM [78] to generate pseudo pixel-level mask annotations and then trained the models with these pseudo pixel-level annotations. Open-vocabulary semantic segmentation [31, 20, 70] aims to perform semantic segmentation on unseen categories. In open-vocabulary semantic segmentation, researchers mainly explored to utilize the pre-trained CLIP [48] to segment unseen categories. Unsupervised segmentation [23, 42, 57] aims to group the pixels belonging to the same objects without using any annotation information.
2.2 Diffusion Models for Segmentation
@@
Diffusion models [16] have exhibited strong image synthesis ability across various text prompts, indicating their strong semantic clue for image segmentation. Inspired by this, researchers started to explore employing the pre-trained diffusion models for segmentation task. For instance, VPD [77] views the pre-trained stable diffusion model as feature extractor, and fine-tunes it for semantic segmentation and depth estimation tasks. TADP [29] introduces image-aligned text prompts to condition the feature extraction of stable diffusion model for semantic segmentation. SLIME [27] adopts stable diffusion model with learnable text embeddings for few-shot segmentation. Additionally, some researchers [24] adopted diffusion framework, in which the model aims to predict the accurate segmentation maps from noisy data using denoising process.
@@
Recently, some researchers explored training-free diffusion models for segmentation. One of the key idea is to employ self-attention map and cross-attention map from pre-trained diffusion model to generate masks of different categories. For instance, DiffuMask [65] first employs diffusion model to generate some synthetic images, and calculates the averaged cross-attention maps as the initial object masks which are refined by AffinityNet. DiffSeg [59] explores grouping self-attention maps at different positions into distinct objects based on the similarities of self-attention maps, followed by a non-maximum suppression operation to assign the pixels to individual objects. OVDiff [26] generates synthetic images by diffusion model, and uses a feature extractor to generate the foreground and background prototypes of different categories for open-vocabulary segmentation. Both DiffSegmenter [63] and T2M [67] refine cross-attention maps of different categories with self-attention maps. To improve cross-attention map, DiffSegmenter and T2M employ additional BLIP [32] and CLIP to carefully design text prompts, which are relatively complex in some degree. Though these aforementioned approaches have exploited self-attention and cross-attention maps for training-free segmentation, we argue that these approaches do not fully utilize the semantic clue existing in self-attention and cross-attention maps. In this paper, we first provide a deep analysis on refining cross-attention map with self-attention map multiple times, and then propose an effective iterative training-free segmentation framework using pre-trained stable diffusion model.
@@ 3 Method
In this section, we first give some preliminaries about stable diffusion model and training-free segmentation. Afterwards, we introduce the motivation of our proposed method. Finally, we describe our proposed iterative refinement framework for training-free segmentation.
@@ 3.1 Preliminaries
Stable diffusion model. Text-to-image (T2I) diffusion model is a type of text-conditioned diffusion models [16], such as DALL-E [49], GLIDE [44], Imagen [55], and stable diffusion [52]. As one of representative T2I models, stable diffusion can generate high-quality images with text prompts, due to the training on large-scale text-image paired dataset (e.g., LAION-5B [56]). It comprises three components: (1) A pre-trained VAE [28] encoder and decoder for image compression and restoration. (2) A text encoder that encodes text descriptions into latent semantic space to condition image generation. (3) A transformer-based denoising UNet [53] that recovers images from noisy data according to the given text descriptions.
Attention modules. Denoising UNet in stable diffusion integrates self-attention and cross-attention mechanisms. The self-attention models the relation between different spatial positions of visual features, while the cross-attention models the relation between visual features and text embeddings. The attention map in both self-attention and cross-attention operations can be calculated by the query and the key , which can be written as
where is normalization factor. Assuming that the visual features have the spatial size of , and the text embeddings have the length of . In self-attention operation, the query and key are both generated by visual features via the linear layers. The generated self-attention map is denoted as , which can reflect spatial similarities of different positions within image. In cross-attention operation, the query and the key are respectively generated by visual features and text embeddings via the linear layers, respectively. The generated cross-attention map is denoted as , which can capture the similarities between image and text.
To achieve training-free segmentation, a straightforward way is setting the text prompts based on given category names, such as ‘A photograph of cat and dog’, then performing the forward process of stable diffusion, and finally generating the corresponding cross-attention map between visual features and text embeddings. Afterwards, we can predict segmentation map by selecting the cross-attention map corresponding to the category name. Instead of directly using the cross-attention map, some researchers [27] proposed to generate refined cross-attention map with self-attention map, which can be written as
@@ 3.2 Motivation
Problem. As mentioned above, self-attention map can be used to refine the cross-attention map. Therefore, a natural question is that, can we improve cross-attention map iteratively with self-attention map, formulated as , where is the index of iterations. We normalize the refined cross-attention map after each iteration. We call this strategy as naive iterative refinement. We initially believe that this naive iterative refinement would yield improved segmentation performance. However, as in Fig. 1(f), it presents performance degradation after several iterations. To better analyze this question, we perform the experiment in Fig. 2. Given the input image (a) and cross-attention map (b), we present the refined cross-attention map at different iterations in (c). It can be observed that the refinement through multiple iterations can not guarantee to generate better refined cross-attention map. We divide the refined results into three situations. The first situation is that the cross-attention map becomes better with the increasing iterations, such as cross-attention map of dog in first row. The second one is the cross-attention map becomes larger, such as the cross-attention map of sheep in the second row. The last one is that the cross-attention map shifts to other regions, exemplified by the cross-attention map of dog showing strong response on the person in third row.
Analysis. Ideally, we argue the refinement with multiple iterations at least presents a stable cross-attention map around corresponding category if the self-attention map is relatively accurate. We further perform a simple experiment by refining cross-attention map with the ground-truth self-attention map in Fig. 2(d). The ground-truth self-attention map is generated by self-correlation of ground-truth masks, which is binary and exclusively has responses within object regions. By using ground-truth self-attention map, the refined cross-attention map becomes stable and accurate after only one iteration. This observation demonstrates that an accurate self-attention map is crucial for accurately refining cross-attention map. As shown in (e), we compare the predicted self-attention map (middle) and ground-truth self-attention map (right) at a given point . Compared to ground-truth self-attention map, predicted self-attention map is relatively noisy and has some weak responses at irrelevant regions. As a result, compared to one iteration, multiple iterations with predicted self-attention map aggregate much irrelevant global information into the refined cross-attention map. Therefore, we argue that it is important to reduce irrelevant global information within self-attention map for iterative refinement. To achieve this goal, we introduce an effective iterative refinement with entropy-reduced self-attention module below.